# Comprehensive survey of hybrid equations of state in neutron star mergers and constraints on the hadron-quark phase transition

### v1

@@ IV Summary and discussion
From our data, we derive an empirical formula for the shift by quantifying $\Delta f_{\mathbf{peak}}$ as a function of $n_{\mathrm{on}}$, $c_{\mathrm{s}}^{2}$ and $\Delta n$ for each set of models with a fixed hadronic part. We demonstrate, how this relation can be used to constrain properties of the hadron-quark phase transition if future GW detections and EoS constraints indicate the presence of such a transition in NS merger remnants. If no clear signs of a transition are present, our relation can also be used to exclude regions of the parameter space and to identify the still viable properties of the transition to deconfined quark matter. This can be complemented by constraints from current heavy-ion experiments such as HADES [95] or future experimental facilities like FAIR [96, 97] and NICA [98, 99].
@@
We have only considered a transitions with a Maxwell construction, a single parametric quark model and three underlying hadronic EoSs. Other types of transitions allowing for coexisting charged phases [15, 60] or a smooth crossover [16, 5] are possible and should also be thoroughly investigated as well as the role of non-convexity of the EoS [100, 101]. It is also important to employ more sophisticated microphysical quark matter models and more hadronic EoSs in future work.
@@ Appendix B EffPT scheme and phase boundaries
Figure 10: (a): Two example phase boundaries for DD2F-based models in the $n$-$T$-plane shown in black and red, respectively. Dashed lines mark the onset of quark deconfinement and solid lines the beginning of the pure deconfined quark matter regime. (b): Same phase boundaries as in (a), but in the $\epsilon_{\mathrm{th}}$-$n$-plane.
@@
Additionally, we slightly modify the effPT scheme for this study compared to Ref. [62]. Due to the ’earlier’ onset of quark deconfinement at finite temperature, a density regime exists where cold matter is in the coexistence phase, but hot material can already be composed of pure deconfined quark matter. Correctly determining the pressure of material in this regime requires knowledge of the pure quark EoS at lower densities. This pressure cannot be directly inferred from a cold, tabulated hybrid EoS, as deconfined quark matter is not present at zero temperature in the required density regime. In Ref. [62], we overcome this problem by using linear extrapolation from the density regime where deconfined quark matter is present at zero temperature to lower densities. Here, we instead infer the pressure of pure quark matter at all densities using Eq. (3) and then add a thermal component with Eq. (7) and $\Gamma_{\mathrm{th}}=4/3$.
@@ Appendix D Estimating fit ranges
As a cross-check we provide an additional estimate of the range of the onset density, where our empirical fit formula (Eq. (2)) is valid. To this end, we infer the maximum onset density $n_{\mathrm{on,max}}$ at which a phase transition will still affect $f_{\mathrm{peak}}$ and beyond which $\Delta f_{\mathrm{peak}}$ frops to zero. This density can be approximated as the maximum density reached in the merger remnant when employing the underlying hadronic EoS model.

### v2

@@ IV Summary and discussion
From our data, we derive an empirical formula for the shift by quantifying $\Delta f_{\mathrm{peak}}$ as a function of $n_{\mathrm{on}}$, $c_{\mathrm{s}}^{2}$ and $\Delta n$ for each set of models with a fixed hadronic part. We demonstrate how this relation can be used to constrain properties of the hadron-quark phase transition if future GW detections and EoS constraints indicate the presence of such a transition in NS merger remnants. If no clear signs of a transition are present, our relation can also be used to exclude regions of the parameter space and to identify the still viable properties of the transition to deconfined quark matter. This can be complemented by constraints from current heavy-ion experiments such as HADES [108] or future experimental facilities like FAIR [109, 110] and NICA [111, 112]. Such complementary information is in particular valuable for small or negligible frequency shifts, which can either imply purely baryonic matter or a transition to quark matter that masquerades baryonic matter [57]. We also note that small frequency shifts of the order of 100 Hz can be caused by specific thermal properties of nucleonic matter [113, 114, 115] or can point to the occurrence of hyperonic degrees of freedom [116, 96].
@@
We have only considered a transitions with a Maxwell construction, a single parametric quark model and three underlying hadronic EoSs. Other types of transitions allowing for coexisting charged phases [15, 62, 55] or a smooth crossover [16, 5] are possible and should also be thoroughly investigated as well as the role of non-convexity of the EoS and magnetic fields [117, 118, 119]. It is also important to employ more sophisticated microphysical quark matter models and more hadronic EoSs in future work.
@@ Appendix B EffPT scheme and phase boundaries
Figure 10: (a): Two example phase boundaries for DD2F-based models in the $n$-$T$-plane shown in black and red, respectively. Dashed lines mark the onset density of quark matter and solid lines the beginning of the pure deconfined quark matter regime. (b): Same phase boundaries as in (a), but in the $\epsilon_{\mathrm{th}}$-$n$-plane.
@@
Additionally, we slightly modify the effPT scheme for this study compared to Ref. [64]. Due to the ’earlier’ appearance of quark matter at finite temperature, a density regime exists where cold matter is in the coexistence phase, but hot material can already be composed of pure deconfined quark matter. Correctly determining the pressure of material in this regime requires knowledge of the pure quark EoS at lower densities. This pressure cannot be directly inferred from a cold, tabulated hybrid EoS, as deconfined quark matter is not present at zero temperature in the required density regime. In Ref. [64], we overcome this problem by using linear extrapolation from the density regime where deconfined quark matter is present at zero temperature to lower densities. Here, we instead infer the pressure of pure quark matter at all densities using Eq. (3) and then add a thermal component with Eq. (7) and $\Gamma_{\mathrm{th}}=4/3$.
@@ Appendix D Estimating fit ranges
As a cross-check we provide an additional estimate of the range of the onset density, where our empirical fit formula (Eq. (2)) is valid. To this end, we infer the maximum onset density $n_{\mathrm{on,max}}$ at which a phase transition will still affect $f_{\mathrm{peak}}$ and beyond which $\Delta f_{\mathrm{peak}}$ drops to zero. This density can be approximated as the maximum density reached in the merger remnant when employing the underlying hadronic EoS model.